The funded status of blank inc.’s defined benefit pension plan and the balances in prior service cost and the net gain-pensions, are displayed ($ = thousands)

2025 beginning balances 2025 ending balances

PBO 2,300 2,501

Plan assets 2,400 2,591

*Funded status 100 90

Prior service cost-AOCI 325 300

Net gain-AOCI 330 300

The retirees were paid $2,700 and the employer contribution to the pension fund was $245,000 at the end of 2025. The expected rate of return on plan assets was 10% and the actuary’s discount rate is 7%. No changes in actuarial estimates and assumptions regarding the PBO.

Determine the following amounts for 2025:

  1. Actual return on plan assets

  2. Loss or gain on plan assets

  3. Service cost

  4. Pension expense

  5. Average remaining service life of active employees (used to determine amortization of net gain)
Answer #1
Answer: 1. Actual return on Plat assets = Ending plan assets - beginning assets - employer contribution + retirees payment = 2,591 - 2,400 - 245 + 270 = $216,000 2. Gain(loss) on plan assets = Actual return - expected return = 216,000 - (10% * 2,400,000-beginning assets) = ($24,000) loss 3. Service cost = Ending Projected benefit obligation - Beginning Projected benefit obligation - Interest cost + Retiree benefits = 2,501 - 2,300 - (7% * 2,300) + 264 = $304,000 4. Pension expense = Interest cost + expected return + Amortization of prior service cost + amortization of net gain + Service cost = Interest cost + expected return + (beginning prior service cost - ending prior cost) + (Beginning net gain - ending net gain - loss on plan asset) + Service cost = (7% * 2,300) + 240 + (325 - 300) + (330 - 300 - 24) + 304 = $736,000 5. Average remaining service life of active employees = (Beginning Net gain - expected return) / Amortization of net gain = (330 - 300) / (330 - 300 - 24) = 5 years
source: Brainly
answered by: Muhammad Anas
Answer #2

Answers:

Actual return on plan assets

$ in '000

Ending Balance

$         2,591

Add Retiree benefits

$             270

less beginning balance

$       (2,400)

less Cash Contributions

$           (245)

Annual Return

$             216

Loss or gain on plan assets

$ in '000

Expected Return

$             240

(2400*10%)

Actual Return

$             216

Loss on plan assets

$               24

Service Cost

$ in '000

Ending Balance of PBO

$         2,501

Add Retiree Benefits

$             270

Less Opening balance of PBO

$       (2,300)

Less Interest Cost

$           (161)

(2300*7%)

Service Cost

$             310

Pension Expense

$ in '000

Service Cost

$             310

add Interest Cost

$             161

(2300*7%)

Add Prior service cost-AOCI Amortization

$               25

(325-300)

Less Expected Return

$           (240)

Less Net gain AOCI Amortization

$               (6)

(330-300-24)

Pension Expense

$             250

Average remaining service life of active employees

$ in '000

Net Gain at beginning

$             330

Less Expected Return

$           (240)

(2400*10%)

Excess net gain over Expected Return

$               90

Amortized net gain per year

$                  6

Average Service period

15

(90/6)